If m^2 = 17 then what is the value of (m - 1)(m + 1) ?
(A) √17 - 1 (B) √17 + 1 (C) 16 (D) 18 (E) 288

Answer by Theo(3458)

m^2 = 17 means that m = +/- sqrt(17)

(m-1)*(m+1) becomes:

(sqrt(17)-1)*(sqrt(17)+1

or:

(-sqrt(17) - 1) * (-sqrt(17) + 1)

if (sqrt(17)-1)*(sqrt(17)+1, then the product of these 2 factors is processed as follows:

sqrt(17) * sqrt(17) = 17
sqrt(17) * 1 = sqrt(17)
-1 * sqrt(17) = -sqrt(17)
-1 * 1 = -1

add these together to get 17 + sqrt(17) - sqrt(17) - 1 = 17 - 1 = 16

if (-sqrt(17)-1)*(-sqrt(17)+1, then the product of these 2 factors is processed as follows:

-sqrt(17) * -sqrt(17) = 17
-sqrt(17) * 1 = -sqrt(17)
-1 * -sqrt(17) = sqrt(17)
-1 * 1 = -1

add these together to get 17 - sqrt(17) + sqrt(17) - 1 = 17 - 1 = 16

answer would be selection C (16).